OpenStudy (anonymous):

If dy/dx = x * square root of (2x^2+1) , and y contains the point (2,4), then the y-intercept of the particular solution is A)0 B) -2/3 C)-1/3 D)-4/3 or E)-1?

OpenStudy (anonymous):

\[d/x=x \sqrt{2x ^{2}+1}\]

OpenStudy (anonymous):

There is the equation in an easier format!

OpenStudy (freckles):

do you know how to integrate?

OpenStudy (anonymous):

Yeah I got the integral of the equation, I'm just not sure how to find the y-intercept!

OpenStudy (freckles):

can i see what you after you integrated

OpenStudy (freckles):

we can find the constant of integration by using the point (2,4) then we can find the y-intercept by putting x=0 and solving for y

OpenStudy (anonymous):

\[1/6(2x^2+1)^{2/3}+c\]

OpenStudy (freckles):

I think you mean the power to be 3/2 right? anyways solve this for C first \[y=\frac{1}{6}(2x^2+1)^\frac{3}{2}+C \\ 4=\frac{1}{6}(2(2)^2+1)^\frac{3}{2}+C\]

OpenStudy (anonymous):

Oops yeah I did!

OpenStudy (anonymous):

I got c=1/2

OpenStudy (anonymous):

oops c=-1/2

OpenStudy (freckles):

\[4=\frac{1}{6}(2(4)+1)^\frac{3}{2}+C \\ 4=\frac{1}{6}(8+1)^\frac{3}{2}+C \\ 4=\frac{1}{6}(9)^\frac{3}{2}+C \\ 4=\frac{1}{6}(3)^3+C \\ 4=\frac{1}{6}(27)+C\] yeah should be -1/2 :)

OpenStudy (freckles):

\[y=\frac{1}{6}(2x^2+1)^\frac{3}{2}-\frac{1}{2}\]

OpenStudy (freckles):

to find the y-intercept set x=0 and solve for y

OpenStudy (anonymous):

I got -1/3! So the answer wold be C

OpenStudy (anonymous):

Can I ask you one more question possibly?

OpenStudy (freckles):

sounds great

OpenStudy (freckles):

sure

OpenStudy (anonymous):

Okay so the question is "Solve each initial value problem." The problem is: \[dy/dx=e ^{x-y}\] and f(0)=2

OpenStudy (freckles):

first you know law of exponents so you know x^(A+B) can be written as x^A times x^B

OpenStudy (freckles):

\[\frac{dy}{dx}=e^{x}e^{-y}\]

OpenStudy (anonymous):

Yep!

OpenStudy (freckles):

no separate your variables

OpenStudy (freckles):

\[e^{y} dy=e^{x} dx\] integrate both sides

OpenStudy (anonymous):

I'm not sure how you integrate with these

OpenStudy (freckles):

\[\int\limits_{}^{}e^u du=e^u+C\]

OpenStudy (freckles):

since d(e^u)/du=e^u

OpenStudy (freckles):

\[\int\limits_{}^{}e^y dy=\int\limits e^x dx \\ e^y+C_1=e^x+C_2 \\ e^y=e^x+C_2-C_1 \\ e^y=e^x+C\]

OpenStudy (freckles):

you can solve for y if you want to

OpenStudy (freckles):

you are given y(0)=2 this means replace x with 0 and y with 2 so you can solve for C

OpenStudy (anonymous):

when i substituted for y and x I got c to be \[e ^{2}-1\]

OpenStudy (freckles):

sounds awesome so pluggin that back into our solution we have \[e^y=e^x+e^2-1\]

OpenStudy (freckles):

you can leave like this but sometimes it is preferred to solve for y

OpenStudy (anonymous):

how would I solve for y?

OpenStudy (freckles):

do you know the inverse function of f(x)=e^x is g(x)=ln(x)

OpenStudy (anonymous):

yep!

OpenStudy (freckles):

take ln( ) of both sides

OpenStudy (freckles):

\[e^y=e^x+e^2-1 \\ \ln(e^y)=\ln(e^x+e^2-1) \\ y \ln(e)=\ln(e^x+e^2-1) \\ y(1)=\ln(e^x+e^2-1) \\ y=\ln(e^x+e^2-1)\]

OpenStudy (anonymous):

Great! Thank you so much! I also had a problem that was finding the initial value and the equation was \[dy/dx=xe ^{y}\] and f(0)=0 and I ended up with \[(x ^{2}*e ^{y})/2 - 1/2\] is that right?

OpenStudy (freckles):

why no equal sign anywhere?

OpenStudy (anonymous):

the second equation is equal to y, just forgot to add it sorry

OpenStudy (freckles):

ok well this is another one you can do by seperation of variables \[\frac{dy}{dx}=xe^y \\ dy=x e^y dx \\ e^{-y} dy=x dx\] and then you integrate both sides

OpenStudy (freckles):

\[\int\limits_{}^{}e^{-y} dy=\int\limits x dx \\ -e^{-y}=\frac{x^2}{2}+C\]

OpenStudy (freckles):

use y(0)=0 to find your constant of integration

OpenStudy (anonymous):

alright i did that so c=1

OpenStudy (freckles):

-1?

OpenStudy (anonymous):

yes!

OpenStudy (freckles):

yep sounds good plug into our solution

OpenStudy (freckles):

\[-e^{-y}=\frac{x^2}{2}-1\] you can leave it as this or solve for y

OpenStudy (freckles):

you can solve for y in a very similar way we did before

OpenStudy (freckles):

your solutions will not always be easy or could be impossible to solve for y

OpenStudy (anonymous):

is this one of those problems?

OpenStudy (freckles):

no it isn't this is is easy :p

OpenStudy (freckles):

you multiply -1 on both sides then take ln( ) of both sides

OpenStudy (freckles):

and one more step remains after

OpenStudy (anonymous):

would it be \[y=\ln ((x ^{2}/2)+1)\]

OpenStudy (freckles):

close but still off

OpenStudy (anonymous):

alright what else do I have to do?

OpenStudy (freckles):

\[-e^{-y}=\frac{x^2}{2}-1 \\ \text{ multiply -1 on both sides } \\ e^{-y}=1-\frac{x^2}{2}\] you cool with that step?

OpenStudy (anonymous):

yep!

OpenStudy (freckles):

take ln( ) of both sides

OpenStudy (freckles):

\[\ln(e^{-y})=\ln(1-\frac{x^2}{2})\]

OpenStudy (freckles):

we can use power rule for logs on the left hand side

OpenStudy (freckles):

that is you can do this: \[-y \ln(e)=\ln(1-\frac{x^2}{2})\]

OpenStudy (freckles):

ln(e)=1 though

OpenStudy (freckles):

\[-y=\ln(1-\frac{x^2}{2})\] the last step that remains is multiplying both sides by -1

OpenStudy (freckles):

\[y=- \ln(1-\frac{x^2}{2})\]

OpenStudy (freckles):

there is something else you can do but we don't have to go too crazy

OpenStudy (freckles):

unless you were trying to match an answer to a choice you had or something

OpenStudy (anonymous):

No I'm not matching it to a choice, what else can we do?

OpenStudy (freckles):

you can use power rule again

OpenStudy (freckles):

\[y=- \ln(\frac{2-x^2}{2}) \\ y=\ln((\frac{2-x^2}{2})^{-1}) \\ y=\ln(\frac{2}{2-x^2})\]

OpenStudy (freckles):

I think that looks pretty but you could also use the quotient rule and write it like this \[y=\ln(2)-\ln(2-x^2)\]

OpenStudy (anonymous):

That does look better!

OpenStudy (anonymous):

Great! Thank you so much for all of your help!!

OpenStudy (freckles):

np :)